How do you solve this problem: x-3y=17
2x+y=17?
What is the distance formula?

Answer by praseenakos@yahoo.com(507)

ANSWER
How do you solve this problem: x-3y=17
2x+y=17?
What is the distance formula?

ANSWER:
x - 3y = 17-----------------------------(1)

2x + y = 17------------------------------(2)


We have different methods to solve such problem.

One of them is elimination method.


For that consider equations 1 and 2


First step is to make coefficient of any of the two variable same.


here we have x and 2x.

so multiply the first equation by 2 so that the coefficient of x becomes 2.


==> 2(x - 3y )= 2*17

Remove the parenthesis,



==> 2*x - 2*3y = 34


==> 2x - 6y = 34----------------------(3)



Now consider equation (2) and (3)




2x - 6y = 34----------------------(3)




2x + y = 17------------------------------(2)



Subtract (2) from (3){ That means subtract left side and right side seperately)



==>2x - 6y - (2x + y) = 34 -17



==>2x - 6y - 2x - y) = 17



==> - 7y = 17


Divide -7 on both sides of the equation



==> - 7y/-7 = 17/7



==> y = 17/7




Now substitute this value in any one of the given equations.


Let's take the second equation,



2x + (- 17/7 )= 17


2x - 17/7 = 17



Add 17/7 from both sides of the equation.



2x -17/7 + 17/7 = 17 + 17/7

2x = 136/7


Divide both sides by 2


==> 2x/2 = 136/2*7


==> x = 68/7



So the required solution is,


x = 68/7 and


y = -17/7


(You can chek this answer by plugging these values in the given equation.)


Hope you understood.



DISTANCE FORMULA:

In algebraic geometry, one can find the distance between two points of the xy-plane using the distance formula. The distance between (x1, y1) and (x2, y2) is given by
d = square root of { (x1 - x2)^2 + ( y1 - y2 ) ^2}

For example,

If A ( 5, 7 ) and B( -1 , 15 )


then by distance formula, the distance betweeen A and B is given by,



d = square root of { (x1 - x2)^2 + ( y1 - y2 ) ^2}


d = square root of { (5 - -1)^2 + ( 7 - 15 ) ^2}


d = square root of { (6)^2 + ( -8 ) ^2}


d = square root of { 36 + 64 }

d = square root of { 100 }

d = 10

That is distance between the points A ( 5, 7 ) and B( -1 , 15 ) is 10 units.


Hope you understood.


Regards,
praseenakos@yahoo.co.in